# Tony Lindeberg

### v6

@@ Abstract
When to apply wavelet analysis to real-time temporal signals, where the future cannot be accessed, it is essential to base all the steps in the signal processing pipeline on computational mechanisms that are truly time-causal.
This paper describes how a time-causal wavelet analysis can be performed based on concepts developed in the area of temporal scale-space theory, originating from a complete classification of temporal smoothing kernels that guarantee non-creation of new structures from finer to coarser temporal scale levels. By necessity, convolution with truncated exponential kernels in cascade constitutes the only permissable class of kernels, as well as their temporal derivatives as a natural complement to fulfil the admissibility conditions of wavelet representations. For a particular way of choosing the time constants in the resulting infinite convolution of truncated exponential kernels, to ensure temporal scale covariance and thus self-similarity over temporal scales, we describe how mother wavelets can be chosen as temporal derivatives of the resulting time-causal limit kernel.
By developing connections between wavelet theory and scale-space theory, we show how the notions of scale normalization in wavelet representations and scale-space representations are closely related regarding scaling properties. We also describe how efficient discrete approximations of the presented theory can be performed in terms of first-order recursive filters coupled in cascade, which obey provably simplifying properties from finer to coarser levels of temporal scales, and thereby guaranteeing well-conditioned numerical implementations. Specifically, we characterize and quantify how the continuous scaling properties transfer to the discrete implementation, demonstrating how the proposed time-causal wavelet representation can reflect the duration of locally dominant temporal structures in the input signals.
We propose that this notion of time-causal wavelet analysis could be a valuable multi-purpose tool for signal processing tasks, where streams of time-dependent signals are to be processed in real time, specifically for signals that may contain local variations over a rich span of temporal scales, or more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
@@ 1 Introduction
Temporal signals can inherently contain different types of structures at different temporal scales. Making such multi-scale relationships in the data explicit, and exploring these relationships, can often substantially improve the performance of signal processing pipelines. For the purpose of handling such scale variations in, not necessarily, only temporal signals, the notion of wavelet analysis has therefore been developed by the signal processing community, to decompose signals into different components at different scales (Ricker 1944, Morlet et al. 1982a; 1982b, Grossmann and Morlet 1984, Hosken 1988, Mallat 1989; 1999, Heil and Walnut 1989, Strang 1989, Daubechies 1992, Meyer 1992, Chui 1992, Ruskain et al. 1992, Donoho and Johnston 1993, Teolis 1998, Walker 1999, Debnath and Shah 2002, Misiti et al. 2007).
@@
While there have been a few wavelet formulations that respect temporal causality (Szu et al. 1992, Vázquez et al. 2005), traditional wavelet theory is, however, largely non-causal, making use of filters that are often symmetric or antisymmetric with respect to the temporal origin. Such an approach is, however, not feasible in real-time situations, when the future cannot be accessed. While one could, in principle, apply non-causal wavelet theory to real-time signals by introducing sufficiently long temporal delays, beyond which the wavelets would be truncated to zero, for time-critical applications, where one wants to minimize the temporal delays, it is of interest to develop time-causal wavelets, that do not make use of any information from the future.
@@
The subject of this paper is to show how a temporal wavelet theory can be developed based on the notion of time-causal temporal scale-space representation, originally developed for video analysis in Lindeberg (2016), and then dedicated for a pure time-causal temporal domain in Lindeberg (2023). This temporal scale-space theory is based on convolution with a special temporal smoothing kernel, denoted the time-causal limit kernel, and which corresponds to the convolution of an infinite set of truncated exponential kernels in cascade, with especially chosen time constants to obtain temporal scale covariance.
@@
Temporal scale covariance in this respect corresponds to self-similarity over temporal scales, so that the temporal scale-space representation computed for a rescaled signal corresponds to a rescaling of the temporal scale-space representation of the original signal, complemented with a shift in the dimension of the temporal scale parameter. By combining this type of time-causal temporal scale-space representation with temporal derivates, we can construct mother wavelets for time-causal wavelet representations from temporal derivatives of the time-causal limit kernel.
@@
Furthermore, by forming the differences between temporal scale-space representations at adjacent temporal scales, we can construct time-causal bandpass wavelet representations, which are formally shown to up to a scaling factor correspond to the time-causal wavelet representation based on first-order temporal derivatives of the time-causal limit kernel. From such a temporal bandpass representation, reconstruction of the original signal is straightforward, by mere addition of the time-causal bandpass representations, which opens up possibilities for real-time manipulation of the signal as decomposed into a time-causal wavelet representation. This situation is structurally very closely related to the corresponding property of the Ricker or Mexican hat wavelet, based on second-order derivatives of the non-causal Gaussian kernel, where reconstruction can be directly performed by integrating the corresponding second-order non-causal wavelet representations over the scale parameter, or the difference-of-Gaussians wavelet, where reconstruction is performed by instead summing the difference-of-Gaussians wavelet representations over temporal scales.
@@
By the time-causal limit kernel constituting the composition of a set of truncated exponential kernels in cascade, the temporal smoothing operation equivalently corresponds to applying a set of first-order integrators in cascade. In terms of discrete implementation, such an operation can be approximated by a set of first-order recursive filters coupled in cascade. Thereby, the resulting implementation is fully time-recursive, meaning that no other temporal memory of the past is needed, beyond the information contained in the temporal scale channels themselves. Specifically, this implies straightforward and very efficient implementation schemes on regular signal processing architectures.
@@
By combining such mother wavelets with different types of normalizations with respect to the scale parameter, either in terms of $L_{p}$-norms or the notion of scale-normalized derivatives, scale selective signal processing algorithms can then be designed, that lead to extrema over scales in the magnitude of the resulting time-causal wavelet representations at temporal scale levels proportional to characteristic durations of different types of temporal structures in the input data (Lindeberg 2017; 2018b; 2018a), with close relationships to the previously studied notion of modulus maxima in wavelet representations (Mallat and Hwang 1992).
@@
In this way, we will demonstrate that significant components in more traditional non-causal wavelet theory can be carried over to a time-causal temporal domain, which opens up for a variety of application areas, where either (i) temporal signals are to be processed in real time, such as for purposes of real-time monitoring, time series analysis and prediction, or as components in control systems, that are to react to a continuous inflow of information, or (ii) analysis of biological or physical signals that needs to be performed in a completely time-causal manner, to be biologically or physically realistic, such as in mathematically based models of biological or physical processes.
@@
Beyond a very few early approaches, this is the first systematic treatment of time-causal wavelet theory, with structurally close relations to non-causal wavelets based on derivatives of the Gaussian kernel. Specifically, we propose that the time-causal wavelet theory to be presented here constitutes a canonical way to define wavelet representations over a time-causal temporal domain, in a corresponding way as non-causal wavelets based on derivatives of the Gaussian kernel, such as the Ricker or Mexican hat wavelets, can be regarded as canonical wavelet representations over a non-causal temporal domain, according to the theory for variation-diminishing convolution transformations.
@@ 1.1 Structure of the presentation
Section 4 then shows how this continuous kernel can be canonically discretized in terms of a set of first-order recursive filters coupled in cascade, complemented with time-causal difference operators to compute discrete approximations of temporal derivatives of the time-causal limit kernel, thereby leading to a family of discrete time-causal wavelet representations as summarized in Section 5.
@@
Section 7 complements with examples of results of computing the proposed time-causal wavelet representations for different types of temporal signals. Finally, Section 8 concludes with a summary and discussion, including relations to related multi-scale signal processing approaches.
@@ 2 Continuous wavelet representations based on temporal derivatives of the time-causal limit kernel
Specifically, since the time-causal limit kernel, that we will use as a fundamental basis for defining the time-causal wavelet representations, does not have any compact closed-form expression, we will in this section devote special emphasis on describing its theoretical properties, so as to be able to infer properties of the derived time-causal wavelet representations, and in this way mediate a better theoretical understanding.
@@ 2.2 The time-causal limit kernel
This time-causal kernel has been derived from criteria concerning the notion of temporal scale space representation, where any continuous signal $f\colon{\mathbb{R}}\rightarrow{\mathbb{R}}$ is to be represented by a temporal scale-space representation $L\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$
@@
over multiple temporal scale levels $\tau\in{\mathbb{R}}_{+}$ in such a way that no new local extrema or zero-crossings may be created from finer to coarser levels of temporal scales. The parameter $c$ is specifically a distribution parameter with $c>1$ that determines the discrete scale levels in units of the variance of the time-causal limit kernel according to [^2] This logarithmic spacing of the temporal scale levels is similar to the spacing generated by a uniform spacing with respect to the canonical logarithmic parameterization of the scale parameter into an effective scale parameter $\tau_{\text{eff}}$ according to $\tau_{\text{eff}}=A+B\log\tau$, as can be derived from axiomatic arguments for a regular spatial scale-space representation (Lindeberg 1993). Compared to a uniform spacing of the scale levels relative to the scale parameter measured in units of the variance $\tau$ of the temporal smoothing kernel, the logarithmic distribution according to (12) also leads to significantly shorter delays, as further described in Lindeberg (2016) Section 4 and Appendix 1. (Lindeberg 2016 Equation (18))
@@ 2.7.2 Scale selection properties for characteristic temporal signals
In relation to normalization of the filters with respect to $L_{p}$-norms, it can specifically be shown that (Lindeberg 1998a Section 9.1) the notion of $\gamma$-normalized derivatives corresponds to normalizing the $n$:th order scale-normalized Gaussian derivative kernels $g_{\zeta^{n}}(t;\;\tau)$ over a one-dimensional domain to constant $L_{p}$-norms over scale
@@
With regard to the above detection of the scale of a 1-D Gaussian blob from extrema in the magnitude of the scale-normalized second-order derivative using the scale normalization power $\gamma_{2}=3/4$, or the detection of the scale of a 1-D Gaussian edge from extrema over scale in the magnitude of the scale-normalized first-order derivative using the scale normalization power $\gamma_{1}=1/2$, it can be noted that both of these scale normalization methods correspond to normalizing the corresponding Gaussian derivative kernels to constant $L_{p}$-norms over scale for the choice of $p=2/3$.
@@ 3.3 Bandpass representations based on the second-order derivatives of the Gaussian kernel
it follows from the fact that the difference between the Gaussian-based temporal scale-space representations at the scales $\tau$ and $\tau+\Delta\tau$ will constitute an approximation of the derivative of the scale-space representation in the scale direction
@@
$L(t;\;\tau_{1})=L(t;\;\tau_{2})+\frac{1}{2}\int_{t=t_{1}}^{t_{2}}(g_{tt}(\cdot;\;\tau)*f(\cdot))(t;\;\tau)\,d\tau.$ (111)
@@ 4.3 Discrete analogue of the time-causal limit kernel
For formulating a discrete analogue of the time-causal limit kernel, one special fact to consider is that while the temporal mean of the first-order recursive filter is structurally similar as for the truncated exponential filter $V=m_{k}$, the expression for the variance of the discrete recursive filter is structurally different
@@ 5 Time-causal wavelet representations for discrete signals
Given that we use the scheme in Algorithm 1 for smoothing a discrete signal stream $f\colon{\mathbb{Z}}\rightarrow{\mathbb{R}}$ with the discrete analogue of the time-causal limit kernel $\Psi_{\text{disc}}(t;\;\tau,c)$, and then compute discrete approximations of temporal derivatives using the time-causal discrete temporal derivative operators $\delta_{t}$ and $\delta_{tt}$ according to (123), complemented with scale normalization according to
@@ 7 Experiments
As can be seen from the results, the dominant responses in the time-causal wavelet representations across scales move from finer to coarser scales with increasing local wavelengths in the signal, in agreement with the desirable properties of a multi-scale wavelet representation, and also in agreement with the scaling properties.
@@
Figure 11 shows corresponding results for an auditory percussion sound of a cymbal (from https://pixabay.com/sound-effects/search/cymbal/ item “Cymbal Crash”). As we can see, the time-causal wavelet transform largely responds over three ranges of temporal scales, with different durations of the responses over these ranges of temporal scales. In this way, the proposed time-causal wavelet transform leads to an intuitive separation of the temporal variations in the signal over different temporal scales.

### v9

@@ Abstract
This paper presents a framework for time-causal wavelet analysis. It targets real-time processing of temporal signals, where data from the future are not available.
The study builds upon temporal scale-space theory, originating from a complete classification of temporal smoothing kernels that guarantee non-creation of new structures from finer to coarser temporal scale levels. We construct temporal wavelets from the temporal derivatives of a special time-causal smoothing kernel, referred to as the time-causal limit kernel, as arising from the classification of variation-diminishing smoothing transformations with the complementary requirement of temporal scale covariance, to guarantee self-similar handling of structures in the input signal at different temporal scales. This enables decomposition of the signal into different components at different scales, while adhering to temporal causality.
The paper establishes theoretical foundations for these time-causal wavelet representations, and maps structural relationships to the non-causal Ricker or Mexican hat wavelets. We also describe how efficient discrete approximations of the presented theory can be performed in terms of first-order recursive filters coupled in cascade, which enables numerically well-conditioned real-time processing with low resource usage. We characterize and quantify how the continuous scaling properties transfer to the discrete implementation, demonstrating how the proposed time-causal wavelet representation can reflect the duration of locally dominant temporal structures in the input signal.
We propose this notion of time-causal wavelet analysis as a generic multi-purpose tool for signal processing tasks, where streams of signals are to be processed in real time, specifically for signals that may contain local variations over a rich span of temporal scales, or more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
@@ 1 Introduction
Temporal signals can inherently contain different types of structures at different temporal scales. Making such multi-scale relationships in the data explicit, and exploring these relationships, can often substantially improve the performance of signal processing pipelines. For the purpose of handling such scale variations in not necessarily only temporal signals, the notion of wavelet analysis has therefore been developed by the signal processing community, to decompose signals into different components at different scales (Ricker 1944, Morlet et al. 1982a; 1982b, Grossmann and Morlet 1984, Hosken 1988, Mallat 1989; 1999, Heil and Walnut 1989, Strang 1989, Daubechies 1992, Meyer 1992, Chui 1992, Ruskain et al. 1992, Donoho and Johnston 1993, Teolis 1998, Walker 1999, Debnath and Shah 2002, Misiti et al. 2007). Due to their scale separating properties and their computational efficiency, wavelet representations are used as a powerful generic building block for representing the information contents in image data and time-dependent signals for a very wide range of signal processing methods and applications.
@@
While there have been a few wavelet formulations that respect temporal causality (Szu et al. 1992, Vázquez et al. 2005), traditional wavelet theory is, however, largely non-causal, making use of filters that are often symmetric or antisymmetric with respect to the temporal origin. Such an approach is, however, not feasible in real-time situations, when the future cannot be accessed. While one could, in principle, apply non-causal wavelet theory to real-time signals by introducing explicit temporal buffers with sufficiently long temporal delays, beyond which the wavelets would be truncated to zero, for time-critical applications, where one wants to minimize the temporal delays, it is of interest to develop truly time-causal wavelets, that do not make use of any information from the future.
@@
The subject of this paper is to show how a principled temporal wavelet theory can be developed based on the notion of time-causal temporal scale-space representation, originally developed for video analysis in Lindeberg (2016), and then dedicated for a pure time-causal temporal domain in Lindeberg (2023). This temporal scale-space theory, which determines canonical filter shapes for the temporal wavelets, is based on convolution with a special temporal smoothing kernel, denoted the time-causal limit kernel, and which corresponds to the convolution of an infinite set of truncated exponential kernels in cascade, with especially chosen time constants to obtain temporal scale covariance.
@@
Temporal scale covariance in this respect corresponds to self-similarity over temporal scales, so that the temporal scale-space representation computed for a rescaled signal corresponds to a rescaling of the temporal scale-space representation of the original signal, complemented with a shift in the dimension of the temporal scale parameter. By combining this type of time-causal temporal scale-space representation with temporal derivates, we can construct mother wavelets for time-causal wavelet representations from temporal derivatives of the time-causal limit kernel. In this way, the resulting time-causal wavelets will have the ability to handle temporal structures at different temporal scales in a provably self-similar manner, and can in this way constitute the foundation for defining time-causal and scale-covariant temporal basis functions, to be used as computational primitives in more composed signal processing or learning-based methods.
@@
Furthermore, by forming the differences between temporal scale-space representations at adjacent temporal scales, we can construct time-causal bandpass wavelet representations, which are here formally shown to up to a scaling factor correspond to the time-causal wavelet representation based on first-order temporal derivatives of the time-causal limit kernel. From such a temporal bandpass representation, reconstruction of the original signal is straightforward, by mere addition over scales of the time-causal bandpass representations, which opens up possibilities for real-time manipulation of the signal as decomposed into a time-causal wavelet representation. This situation is structurally very closely related to the corresponding property of the Ricker or Mexican hat wavelet, based on second-order derivatives of the non-causal Gaussian kernel, where reconstruction, as will be shown here, can be directly performed by integrating the corresponding second-order non-causal wavelet representations over the scale parameter, or the difference-of-Gaussians wavelet, where reconstruction is performed by instead summing up the difference-of-Gaussians wavelet representations over the temporal scales.
@@
By the time-causal limit kernel constituting the composition of a set of truncated exponential kernels in cascade, the temporal smoothing operation equivalently corresponds to applying a set of first-order integrators in cascade. In terms of discrete implementation, such an operation can be well approximated by a set of first-order recursive filters coupled in cascade. Thereby, the resulting implementation is fully time-recursive, meaning that no other temporal memory of the past is needed, beyond the information contained in the temporal scale channels themselves. Specifically, this implies straightforward and very efficient implementation schemes on regular signal processing architectures.
@@
By combining such mother wavelets with different types of normalizations with respect to the scale parameter, either in terms of $L_{p}$-norms or the notion of scale-normalized derivatives, scale selective signal processing algorithms can then be designed, that lead to extrema over scales in the magnitude of the resulting time-causal wavelet representations at temporal scale levels proportional to characteristic durations of different types of temporal structures in the input data (Lindeberg 2017; 2018b; 2018a), with close relationships to the previously studied notion of modulus maxima in non-causal wavelet representations (Mallat and Hwang 1992).
@@
In this way, we will demonstrate that significant components in more traditional non-causal wavelet theory can be carried over to a time-causal temporal domain, which opens up for a variety of application areas, where either (i) temporal signals are to be processed in real time, such as for purposes of real-time monitoring, time series analysis and prediction, including learning-based systems as detailed further in Section 8, or as components in control-loop systems, that are to react to a continuous inflow of information, as well as (ii) analysis of biological or physical signals that needs to be performed in a completely time-causal manner, to be biologically or physically realistic, such as in mathematically based models of biological or physical processes.
@@
Beyond a very few early approaches, this is the first systematic treatment of time-causal wavelet theory, with structurally close relations to non-causal wavelets based on derivatives of the Gaussian kernel. Specifically, we propose that the time-causal wavelet theory to be presented here constitutes a canonical way to define wavelet representations over a time-causal temporal domain, in a corresponding way as non-causal wavelets based on derivatives of the Gaussian kernel, such as the Ricker or Mexican hat wavelets, can be regarded as canonical wavelet representations over a non-causal temporal domain, according to the theory for variation-diminishing convolution transformations to be described in Section 2.3.
@@ 1.1 Structure of the presentation
Section 4 then shows how this continuous kernel can be canonically discretized in terms of a set of first-order recursive filters coupled in cascade, complemented with time-causal difference operators, to compute discrete approximations of temporal derivatives of the time-causal limit kernel, thereby leading to a family of discrete time-causal wavelet representations as summarized in Section 5.
@@
Section 7 complements with examples of results of computing the proposed time-causal wavelet representations for different types of temporal signals. Finally, Section 8 concludes with a summary and discussion, including relations to related multi-scale signal processing and learning approaches.
@@ 2 Continuous wavelet representations based on temporal derivatives of the time-causal limit kernel
The intention behind this treatment is to lay out the theoretical foundations for using the proposed time-causal wavelets as computational primitives in more composed signal processing and learning approaches. Specifically, since the time-causal limit kernel, that we will use as a fundamental basis for defining the time-causal wavelet representations in terms of derivatives thereof, does not have any compact closed-form expression, we will in this section devote special emphasis on describing its theoretical properties, so as to be able to infer properties of the derived time-causal wavelet representations, and in this way mediate a better theoretical understanding.
An effectively more compact description of some of the main concepts and results will then be given in Section 3.
@@ 2.2 The time-causal limit kernel
This time-causal kernel has been derived from criteria concerning the notion of temporal scale space representation, where any continuous signal $f\colon{\mathbb{R}}\rightarrow{\mathbb{R}}$ is to be represented by a temporal scale-space representation $L\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ (see Lindeberg 2023 Equation (26))
@@
over multiple temporal scale levels $\tau\in{\mathbb{R}}_{+}$ in such a way that no new local extrema or zero-crossings may be created from finer to coarser levels of temporal scales. The parameter $c$ is specifically a distribution parameter with $c>1$, that determines the discrete scale levels in units of the variance of the time-causal limit kernel according to [^2] This logarithmic spacing of the temporal scale levels is similar to the spacing generated by a uniform spacing with respect to the canonical logarithmic parameterization of the scale parameter into an effective scale parameter $\tau_{\text{eff}}$ according to $\tau_{\text{eff}}=A+B\log\tau$, as can be derived from axiomatic arguments for a regular spatial scale-space representation (Lindeberg 1993). Compared to a uniform spacing of the scale levels relative to the scale parameter measured in units of the variance $\tau$ of the temporal smoothing kernel, the logarithmic distribution according to (12) also leads to significantly shorter delays, as further described in Lindeberg (2016) Section 4 and Appendix 1. (Lindeberg 2016 Equation (18))
@@ 2.7.2 Scale selection properties for characteristic temporal signals
In relation to normalization of the filters with respect to $L_{p}$-norms, it can specifically be shown (Lindeberg 1998a Section 9.1) that the notion of $\gamma$-normalized derivatives corresponds to normalizing the $n$:th order scale-normalized Gaussian derivative kernels $g_{\zeta^{n}}(t;\;\tau)$ over a one-dimensional domain to constant $L_{p}$-norms over scale
@@
With regard to the above detection of the scale of a 1-D Gaussian blob from extrema over scale in the magnitude of the scale-normalized second-order derivative using the scale normalization power $\gamma_{2}=3/4$, or the detection of the scale of a 1-D Gaussian edge from extrema over scale in the magnitude of the scale-normalized first-order derivative using the scale normalization power $\gamma_{1}=1/2$, it can be noted that both of these scale normalization methods correspond to normalizing the corresponding Gaussian derivative kernels to constant $L_{p}$-norms over scale for the choice of $p=2/3$.
@@ 3.3 Bandpass representations based on the second-order derivatives of the Gaussian kernel
it follows from the fact that the difference between the non-causal Gaussian-based temporal scale-space representations at the scales $\tau$ and $\tau+\Delta\tau$ will constitute an approximation of the derivative of the scale-space representation in the scale direction
@@
$L(t;\;\tau_{1})=L(t;\;\tau_{2})-\frac{1}{2}\int_{t=t_{1}}^{t_{2}}(g_{tt}(\cdot;\;\tau)*f(\cdot))(t;\;\tau)\,d\tau.$ (111)
@@ 4.3 Discrete analogue of the time-causal limit kernel
For formulating a discrete analogue of the time-causal limit kernel, one special fact to consider is that while the temporal mean $M$ of the first-order recursive filter is structurally similar as for the truncated exponential filter $M=\mu_{k}$, the expression for the variance of the discrete recursive filter is structurally different
@@ 5 Time-causal wavelet representations for discrete signals
Given that we use the scheme in Algorithm 1 for smoothing a discrete signal stream $f\colon{\mathbb{Z}}\rightarrow{\mathbb{R}}$ with an approximation of the discrete analogue of the time-causal limit kernel $\Psi_{\text{disc}}(t;\;\tau,c)$, and then compute discrete approximations of temporal derivatives using the time-causal discrete temporal derivative operators $\delta_{t}$ and $\delta_{tt}$ according to (123), complemented with scale normalization according to
@@ 7 Experiments
As can be seen from the results, the dominant responses in the time-causal wavelet representations across scales move from finer to coarser scales with increasing local wavelengths in the signal, in agreement with the desirable properties of a multi-scale wavelet representation, and also in agreement with the scaling properties. Due to the exponential increase of the local wavelengths in the input signal as function of time, the corresponding temporal scale estimates ought to also increase exponentially over time. With the logarithmic parameterization of the temporal scale parameter used in the figures, we experimentally obtain scale dependencies that are well described as linear on the logarithmic scale, thus with a very good match between the continuous theory and the discrete implementation.
@@
Figure 11 shows corresponding results for an auditory percussion sound of a cymbal (from https://pixabay.com/sound-effects/search/cymbal/ item “Cymbal Crash”). As we can see, the time-causal wavelet transform largely responds over three ranges of temporal scales, with different durations of the responses over these ranges of temporal scales, and as corresponding to vibrations in the cymbal over different ranges of frequencies. In this way, the proposed time-causal wavelet transform leads to an intuitive separation of the temporal variations in the signal over different temporal scales.